A 0.050 mg dose of iodine-131 was administered to a patient to treat hyperthyroidism. What is the remaining mass after 32.4 days? The half-life of iodine-131 is 8.1 days.
Answer #1

11) t = 32.4 days

t1/2 = 8.1 days

A = 0.050 mg

A0 = ?

t= ln( A0/A) X  t1/2/0.693

32.4 = ln(A0/0.05) x 8.1/0.693

2.772 = ln(A0) - ln(0.05)

-0.224 = ln(A0)

A0 = 0.597 mg